# MAnycast Reloaded: a Tool for an Open, Fast, Responsible and Efficient Daily Anycast Census

### v1

@@ Abstract.
IP anycast is a widely adopted technique in which an address is replicated at multiple locations, to, e.g., reduce latency and enhance resilience. Due to anycast’s crucial role on the modern Internet, earlier research introduced tools to perform anycast censuses. The first, iGreedy, uses latency measurements from geographically dispersed locations to map anycast deployments. The second, MAnycast2, uses anycast to perform a census of other anycast networks. MAnycast2’s advantage is speed, performing an Internet-wide census in 3 hours, but it suffers from problems with accuracy and precision. Inversely, iGreedy is highly accurate but much slower. On top of that, iGreedy has a much higher probing cost.
In this paper we address the shortcomings of both systems and present MAnycast Reloaded (MAnycastR). Taking MAnycast2 as a basis, we completely redesign its measurement pipeline, and add support for distributed probing, additional protocols (UDP, TCP and IPv6) and latency measurements similar to iGreedy. We validate MAnycastR on an anycast testbed with 32 globally distributed nodes, compare against an external anycast production deployment and extensive latency measurements with RIPE Atlas, and cross-check over 60% of detected anycast prefixes against operator ground truth. This shows that MAnycastR achieves high accuracy and precision. We make continual daily MAnycastR censuses available to the community and release the source code of the tool under a permissive open source license.
copyright: none
@@ 1. Introduction
Anycast is a technique where an IP address is made available in multiple, autonomous locations such that packets are routed to one of several available locations (Lindqvist and Abley, 2006). This technique is widely used to satisfy client requests from multiple locations to ensure low-latency responses. Anycast also allows for load-balancing by distributing traffic over multiple sites. Finally, it can enhance resilience by ensuring the availability of a service in multiple locations that can absorb traffic in case one location goes down. Prominent examples of services that use anycast include the Domain Name System (DNS) (Sommese et al., 2021) and Content Delivery Networks (CDNs) (Calder et al., 2015; Koch et al., 2021).
@@
Given its widespread use, determining how widely and for what services anycast is used is vital to our understanding of the modern-day Internet. Several techniques have been developed to perform Internet-wide IP anycast censuses. The two most prominent ones are iGreedy (Cicalese et al., 2015) and MAnycast2 (Sommese et al., 2020). The first, iGreedy, uses latency measurements from geographically dispersed locations to infer anycast use. iGreedy detects anycasted prefixes by looking for “speed-of-light violations” (i.e., latencies from two locations to a target that are physically impossible). A full census of the IPv4 address space using iGreedy takes days (Cicalese et al., 2015). The second technique, MAnycast2, involves probing from multiple Vantage Points (VPs) using an anycast source address. A target will send its responses to the VP that is closest in terms of BGP routing. A unicast target will therefore send responses from one location to a single VP, whereas an anycast target’s responses are sent from multiple locations to multiple VPs. The advantages of MAnycast2 are its speed, performing a full census of the routable IPv4 space in under 3 hours, and that it needs a far more modest probing budget to discover anycasted prefixes. Unfortunately, MAnycast2 also suffers from problems with accuracy and precision. It is particularly prone to misclassifying targets as anycast for cases where responses to probes end up at a low number of VPs (Sommese et al., 2020).
In this work, we tackle these shortcomings and present MAnycast Reloaded (MAnycastR). Taking MAnycast2 as a starting point, we design and build a new measurement pipeline from the ground up. We improve accuracy by adding support for synchronized probing. We also incorporate the latency-based measurement algorithm from iGreedy. MAnycastR’s streamlined implementation outperforms MAnycast2 in terms of speed, allowing for a census and accompanying latency-based validation to be performed daily. In other words: MAnycastR is faster than MAnycast2 (and thus much faster than iGreedy), while being just as accurate as iGreedy.
The contributions of this paper are that we:
  • Design and implement MAnycastR from the ground up;
  • Add support for IPv6, UDP- and TCP-based probing;
@@
  • Validate MAnycastR’s performance against ground truth from operators, measurements on an anycast production deployment, commercial datasets, and extensive latency-based measurements through RIPE Atlas;
  • Release daily anycast censuses for the full IPv4 address space and state-of-the-art IPv6 hitlists [^1] https://github.com/anycast-census/anycast-census;
  • Provide a user-friendly dashboard to query the census [^2] https://manycast.net;
  • Release the full source code for MAnycastR under a permissive open source license (upon publication).
@@
The remainder of this paper is organized as follows. First, we provide background on iGreedy and MAnycast2 and discuss other related work on anycast measurement methodologies in §2. In §3 we discuss the requirements for MAnycastR. Next, in §4 we introduce the design and implementation of MAnycastR. §5 discusses our extensive validation. Finally, we conclude in §6.
Figure 1. Using latency measurements from VPs (green, purple) for GCD-based anycast site (black) detection.
@@ 2. Background and Related Work
This section provides background information on iGreedy and its latency-based anycast detection method and on the MAnycast2 anycast-based approach. We end with a discussion of other related work on anycast detection approaches.
@@
The first comprehensive anycast census technique is iGreedy, introduced by Cicalese et al. in 2015 (Cicalese et al., 2015). iGreedy’s anycast detection approach is based on using the Great-Circle Distance (GCD). iGreedy leverages assumptions about the speed of light to determine the maximum geographical distance an Internet packet travels based on its Round-Trip Time (RTT).
By capturing latency data from multiple VPs to a probed target, iGreedy determines the area in which the target must reside using GCD. Figure 1 illustrates this principle. A probe is sent from each VP (green and purple dots). The RTT for these probes is then used to draw a circle indicating the distance a packet can travel. The point where the circles intersect is then used to approximate the location of the target. Wherever there are non-overlapping circles (green and purple circles in the example), iGreedy infers the target to be anycast as it must be at multiple locations. Enumeration is achieved by determining the minimum set of areas required to satisfy a valid location for all VPs, thereby not violating any speed-of-light constraints. Next, geolocation is achieved by picking the most likely city (based on population counts) where the anycast site resides in each area.
The GCD approach relies on assumptions about the speed at which packets travel. iGreedy’s default is the speed of light in fibre ($\sim$200,000 km/s). This estimate discards, e.g., delays due to buffering. Therefore, the GCD approach likely underestimates both the number of anycasted IP prefixes and the number of sites per prefix. In particular, it may fail to detect anycast deployed in close proximity (regional anycast).
Figure 2. Using an anycast setup (light blue) to detect anycast. Probes to a unicast site (orange) result in responses to a single VP (blue), probes to anycast sites (green) result in responses to multiple VPs (red).
@@ 2.2. MAnycast2
MAnycast2 takes a different approach by using anycast to detect anycast. The basic principle behind MAnycast2 is based on an intuition by De Vries et al. (de Vries et al., 2020) who, in their paper, test “Verfploeter”, an approach to map anycast catchments, with a catchment defined as all prefixes that send their traffic to a given point of presence of an anycast service. When De Vries et al. tested catchment stability, they noticed that some prefixes showed up in the catchment of many locations. They intuited that these prefixes were themselves using anycast, and confirmed some of this with ground truth about DNS root servers.
MAnycast2 leverages this to map anycast. Figure 2 illustrates the techique, showing how probing a unicast address from an anycast setup leads to responses being received by a single VP, whereas probing an anycast address leads to responses at multiple VPs. As such, if a single VP receives all responses for a given target, MAnycast2 infers it to be unicast, and when multiple VPs receive responses, the target is inferred to be anycast. However, there are corner cases where this intuition fails, resulting in the following possibilities: True positive (TP) - an anycast target has sites respond to different VPs; True negative (TN) - a unicast target consistently responds to the same VP; False positive (FP) - a unicast target responds to multiple VPs; False negative (FN) - an anycast target has all sites respond to the same VP.
The errors in this approach can occur for the following reasons: A FP can occur when the unicast target is ‘in between’ two VPs, meaning there are equal-cost (in terms of BGP hops) paths to the VPs where, e.g., Equal-Cost Multi-Path (ECMP) routers send responses to both sites. Furthermore, route flips may cause a target to respond to a different VP when a flip occurs between sequential MAnycast2 probes. A FN can occur when a probed target’s anycast sites are in the catchment of a single VP, meaning all sites of the target will respond to the same VP. To minimize FNs the measuring anycast infrastructure must consist of many geographically dispersed VPs, lowering the odds of a probed anycast deployment being entirely in the catchment of a single VP.
Whilst MAnycast2 is able to detect and enumerate anycast, iGreedy detects anycast with fewer FPs, and is able to geolocate anycast sites at a city-level granularity. At the same time, iGreedy suffers FNs when sites are geographically close. To find a middle ground between probing cost and reducing FPs, the MAnycast2 paper proposes a combined approach, in which an anycast-based measurement creates a target set of potential anycast prefixes, which is then refined using an iGreedy measurement using a much larger number of VPs. In this paper, we build on this combined approach and publicly share daily census results of both methodologies.
@@ 2.3. Other related work
Anycast is used extensively for the DNS, particularly for the DNS root servers. For troubleshooting purposes, it is useful to identify which instance a client reaches when it sends a query. For this reason, RFC 4892 specifies a mechanism for identifying anycast sites relying on CHAOS-class DNS records, where each site discloses its identity in a TXT record (Woolf and Conrad, 2007). Through a CHAOS query, a client can then determine which site it reaches. CHAOS queries have been used to study anycast deployments. In 2011 Fan et al. proposed “ACE” (Anycast Characterization and Evaluation) that uses CHAOS records in combination with traceroute to detect and enumerate anycast (Fan et al., 2013). Two years later Fan et al. improved the scope of this approach by proposing Internet-class (IN) DNS records that can be requested using existing recursive DNS infrastructure (Fan et al., 2011). They find that 10 k VPs are required for a recall of 80% when enumerating the number of anycasted replicas. Fan et al.’s approach is only usable for the DNS, whereas iGreedy and MAnycast2 are protocol-agnostic.
Apart from iGreedy and MAnycast2, Bian et al. introduced one other approach for service-agnostic anycast detection. They passively collected BGP announcements to detect anycast by looking for geographically diverse upstreams (Bian et al., 2019). However, their methodology suffered from false positives due to remote peering, since remote peering allows a unicast prefix to have upstreams in geographically distant locations.
In addition to developing iGreedy, Cicalese et al. also performed a finite longitudinal census of IP anycast on a monthly basis with iGreedy (covering December 2015 to May 2017) (Cicalese and Rossi, 2018). In this work, we significantly expand on this by providing a fine-grained daily longitudinal dataset that will enable detailed study of the IP anycast landscape.
@@ 3. System requirements
We set the following goals in this work. We want to build a system that: i) combines anycast- and latency-based GCD measurements; ii) is accurate and precise; and iii) performs a responsible measurement in terms of probing budget. Furthermore, we want to release a continual daily census for the community and release the toolchain as open source for others to build on.
To achieve these goals we set the following requirements:
@@
    Accuracy - MAnycastR must minimise false positives and false negatives. Additionally, MAnycastR must convey confidence in results through independently listing the classification (unicast, anycast, or unresponsive) for the anycast-based and GCD approach.
@@
    Precision - The differences between subsequent censuses due to measurement errors must be minimised. In other words: MAnycastR must be precise over time.
@@
    Responsibility - MAnycastR must perform measurements with low impact on the platform it runs on and the Internet as a whole. MAnycastR must support low probing rates without impacting accuracy and precision.
@@
    Increased coverage - iGreedy and MAnycast2 only support ICMPv4 probing. We want our system to also support IPv6, transport layer protocols (TCP, UDP) and service-aware probing.
@@
    Robustness - A daily census requires a robust measurement system. External issues must be handled properly and be contained, e.g., a VP outage must not impact ongoing measurements besides that particular site.
@@
    Timeliness - A full IPv4 census at /24-prefix granularity, and a hitlist-based IPv6 census must complete fast enough to be able to perform a census for multiple protocols in a single day.
  7. R7
    Ease of deployment - It must be easy to deploy MAnycastR on different anycast infrastructures that may exist for diverse hosting providers.
  8. R8
    Security - To protect measurement integrity, we must secure communication between components.
  9. R9
    Replicability - Independent parties must be able to run MAnycastR and make changes to the code.
  10. R10
    Efficiency - MAnycastR must support high probing rates and deployment on hosts with limited resources.
  11. R11
    Scalability - MAnycastR must function on anycast deployments of all sizes.
@@ 4. Design and Measurement Setup
In this section, we detail the design of the MAnycastR system and show how it meets the requirements from §3. We then explain how we deploy MAnycastR for validation and in production to create daily anycast censuses.
4.1. System design
In this work we reimplement, from scratch, the MAnycast2 approach pioneered by Sommese et al. (see §2.2). Furthermore, we integrate a latency-based approach inspired by iGreedy, and include an improved version of the iGreedy iterative enumeration and geodetection process that significantly reduces processing time, from hours to minutes.

### v3

# LACeS: An Open, Fast, Responsible, and Efficient Longitudinal Anycast Census System

@@ Abstract.
IP anycast replicates an address at multiple locations to reduce latency and enhance resilience. Due to anycast’s crucial role in the modern Internet, earlier research introduced tools to perform anycast censuses. The first, iGreedy, uses latency measurements from geographically dispersed locations to map anycast deployments. The second, MAnycast2, uses anycast to perform a census of other anycast networks. MAnycast2’s advantage is speed and coverage but suffers from problems with accuracy, while iGreedy is highly accurate but slower using author-defined probing rates and costlier.
In this paper we address the shortcomings of both systems and present LACeS (Longitudinal Anycast Census System). Taking MAnycast2 as a basis, we completely redesign its measurement pipeline, and add support for distributed probing, additional protocols (DNS over UDP, TCP SYN/ACK, and IPv6) and latency measurements similar to iGreedy. We validate LACeS on an anycast testbed with 32 globally distributed nodes, compare against an external anycast production deployment, extensive latency measurements with RIPE Atlas and cross-check over 60% of detected anycast using operator ground truth that shows LACeS achieves high accuracy. Finally, we provide a longitudinal analysis of anycast, covering 17+months, showing LACeS achieves high precision. We make continual daily LACeS censuses available to the community and release the source code of the tool under a permissive open source license.
Internet Measurement, Anycast, Internet Topology, Routing, IP
journalyear: 2025copyright: ccconference: Proceedings of the 2025 ACM Internet Measurement Conference; October 28–31, 2025; Madison, WI, USAbooktitle: Proceedings of the 2025 ACM Internet Measurement Conference (IMC ’25), October 28–31, 2025, Madison, WI, USAdoi: 10.1145/3730567.3764484isbn: 979-8-4007-1860-1/2025/10ccs: Networks Network measurement
@@ 1. Introduction
Anycast is a technique where an IP address is made available in multiple, independent sites such that packets are routed to one of several available sites (Lindqvist and Abley, 2006). This technique is widely used to satisfy client requests from multiple sites to ensure low-latency responses. Anycast also allows for load-balancing by distributing traffic over multiple sites. Finally, it can enhance resilience by ensuring the availability of a service in multiple sites that can absorb traffic in case one site goes down. Prominent examples of services that use anycast include the Domain Name System (DNS) (Sommese et al., 2021) and Content Delivery Networks (CDNs) (Calder et al., 2015; Koch et al., 2021).
@@
Given its widespread use, determining how widely and for what services anycast is used is vital to our understanding of the modern-day Internet. Several techniques have been developed to perform Internet-wide IP anycast censuses. The two most prominent are iGreedy (Cicalese et al., 2015 (Cicalese et al., 2015)) and MAnycast2 (Sommese et al., 2020 (Sommese et al., 2020)). The first, iGreedy, uses latency measurements from geographically dispersed Vantage Points (VPs) locations to infer anycast use. iGreedy detects anycasted prefixes by looking for “speed-of-light violations” – latencies from multiple VPs to a target that are physically impossible if the target is in a single location. A census of the ping-responsive IPv4 address space using iGreedy takes days (Cicalese et al., 2015) on currently available community measurement platforms, which limit packet rate, and critically relies on the VPs having an accurate location. The second technique, MAnycast2 (Sommese et al., 2020), involves probing from multiple VPs using an anycast source address. A target will send its responses to the VP that is closest in terms of BGP routing. A unicast target will therefore send responses from one site to a single VP, whereas an anycast target’s responses can be sent from multiple sites to multiple VPs. Because networks that can support anycast experiments are typically well-provisioned, MAnycast2 can infer which IPv4 ping-responsive prefixes are anycast in under 3 hours. Unfortunately, MAnycast2 cannot infer where anycast sites are, and can misclassify a prefix as anycast if routing from a target changes during the measurement.
In preliminary work evaluating MAnycast2, Sommese et al. suggested a future pipeline where MAnycast2 identifies candidate anycast prefixes for followup probing with iGreedy (Sommese et al., 2020). In this work, we design and build a Longitudinal Anycast Census System (LACeS) inspired by that model, that improves MAnycast2’s accuracy through synchronized probing, and incorporates the latency-based measurement algorithm from iGreedy. LACeS’s streamlined implementation outperforms MAnycast2 in terms of speed, allowing for a census and accompanying latency-based validation to be performed daily. The contributions of this paper are that we:
  • Re-engineer MAnycast2 and iGreedy to improve accuracy, coverage (IPv6, UDP- and TCP-based probing) and more;
@@
  • Validate LACeS’s performance against operator ground truth, measurements on an anycast production deployment, commercial datasets, and latency measurements through RIPE Atlas;
  • Release daily anycast censuses for the ping-responsive IPv4 address space and state-of-the-art IPv6 hitlists beginning Mar’24 (git, 2025), accessible with a user-friendly dashboard (LACeS Project, 2025);
  • Release the full source code for LACeS (Hendriks, 2025a).
@@
The remainder of this paper is organized as follows. First, we provide background on iGreedy and MAnycast2 and discuss related work in §2. In §3 we discuss the requirements for LACeS. Next, in §4 we introduce the design and implementation of LACeS. §5 discusses our extensive performance evaluation followed with ground truth validation in §6. Next, we provide a longitudinal analysis in §7 and discuss LACeS and future steps in §8. Finally, we conclude in §9.
Figure 1. Using latency measurements from four VPs (SLC, PHX, BNA, CHS) for GCD-based anycast site (LAS, ATL) detection. These latency measurements imply at least two sites. iGreedy infers the location of each site using the highest populated city in the intersection.
@@ 2. Background and Related Work
This section provides background information on iGreedy and MAnycast2. We end with a discussion of other related work.
@@
iGreedy, introduced by Cicalese et al. in 2015 (Cicalese et al., 2015), uses the Great-Circle Distance (GCD) to determine the maximum geographical distance an Internet packet travels based on its Round-Trip Time (RTT). Figure 1 illustrates this principle – by capturing latency data from multiple VPs to a probed target, iGreedy determines the areas in which a target could reside. When there are non-overlapping circles (green and purple circles in Figure 1), iGreedy infers the target to be anycast. iGreedy enumerates anycast sites by determining the minimum set of areas and infers the geolocation by selecting the highest populated city in each area.
The GCD approach relies on assumptions about the speed at which packets travel. iGreedy’s default is the speed of light in fibre ($\sim$200,000 km/s). This estimate discards, e.g., delays due to buffering. Therefore, the GCD approach can underestimate both the number of anycasted IP prefixes and the number of sites per prefix. In particular, it may fail to detect anycast deployed in close proximity.
Figure 2. Using an anycast setup (light blue nodes – SLC, PHX, BNA, CHS) to detect anycast. Probes to a unicast site (orange lines to OKC) results in responses to a single VP (BNA), while probes to anycast sites (green lines to LAS and ATL) result in responses to multiple VPs (red lines to SLC, CHS).
@@ 2.2. MAnycast2
MAnycast2 (Sommese et al., 2020) is based on an intuition by De Vries et al. (de Vries et al., 2017). De Vries et al. developed “Verfploeter” to map anycast catchments, defined as the set of prefixes that send their traffic to a given site of an anycast service. They inferred catchments by pinging responsive addresses – a passive VP – in each prefix using an anycast source address, noting the site that responses arrive at. When De Vries et al. tested catchment stability, they noticed that some passive VPs were in the catchment of many sites. They intuited that some of the prefixes associated with these passive VPs were themselves anycast, and confirmed this with ground truth.
Figure 2 illustrates how MAnycast2 leverages anycasted VPs to detect anycast. Probing an anycast address can lead to responses at multiple anycasted VPs – a true positive (TP). Probing a unicast address from multiple anycasted VPs can lead to responses being received by a single VP – a true negative (TN).
However, there are cases where this intuition fails. A false positive (FP) occurs when there are equal-cost BGP paths to multiple anycasted VPs, so Equal-Cost Multi-Path (ECMP) routers can send responses to multiple sites. Furthermore, routing dynamics may cause a unicast target to respond to a different anycast VP when a route changes between sequential MAnycast2 probes. A false negative (FN) occurs when a probed target’s anycast sites are in the catchment of a single anycasted VP, meaning all sites of the target will respond to the same VP. To minimize FNs the measuring anycast infrastructure must consist of many topologically dispersed VPs, lowering the odds of a probed anycast deployment being entirely in the catchment of a single VP.
Whilst MAnycast2 is able to detect anycast using few VPs, iGreedy detects anycast with fewer FPs (Sommese et al., 2020), and is able to geolocate anycast sites at a city-level granularity (Cicalese et al., 2015). At the same time, iGreedy suffers FNs when sites are geographically close. To find a middle ground between probing cost and reducing FPs, Sommese et al. suggested a future pipeline where MAnycast2 identifies candidate anycast prefixes for followup probing with iGreedy using a much larger number of VPs (Sommese et al., 2020). In this paper, we build this pipeline using both methodologies and publicly share daily census results.
@@ 2.3. Other related work
Anycast measurement methods – Anycast is used extensively for the DNS, particularly for the DNS root servers. Because it is useful to identify which site a client uses for troubleshooting purposes, RFC 4892 specifies a mechanism for sites to disclose their identity in a CHAOS-class DNS TXT record (Woolf and Conrad, 2007). DNS server operators typically encode a location string in their identity strings, so CHAOS TXT queries are useful when studying anycast deployments. In 2011, Fan et al. proposed “ACE” (Anycast Characterization and Evaluation) that uses CHAOS records in combination with traceroute to detect and enumerate anycast (Fan et al., 2011). In 2013, Fan et al. improved the scope of this approach by proposing Internet-class (IN) DNS records that can be requested using existing recursive DNS infrastructure (Fan et al., 2013). They reported that 10 k VPs were required for a recall of 80% when enumerating the number of anycasted replicas. Because it relies on CHAOS queries, Fan et al.’s approach is DNS-specific.
In 2019, Bian et al. passively collected BGP announcements to detect anycast by looking for geographically diverse upstreams (Bian et al., 2019). Their method is service-agnostic, as it does not rely on DNS records. However, their methodology suffered from false positives due to remote peering, because remote peering allows a unicast prefix to have upstreams in geographically distant locations.
Anycast censuses – In addition to developing iGreedy, Cicalese et al. performed monthly iGreedy censuses of anycast between December 2015 and May 2017 (Cicalese and Rossi, 2018). In this work, we expand on this by providing a fine-grained daily longitudinal dataset that will enable detailed study of the IP anycast landscape.
Current censuses of anycast are largely commercial. BGPTools, a popular platform amongst operators, offers a public list of anycast prefixes using a methodology similar to MAnycast2 (BGPTools, 2025) with four VPs as of Sep. ’25. Geolocation database providers like IPInfo (ipinfo.io, 2025) have a latency-based anycast classifications in their dataset. Whereas these censuses only report anycast prefixes, we additionally enumerate and geolocate anycast. We compare results with BGPTools and IPInfo in §5.8.
@@ 3. System requirements
MAnycast2 suffers from a high false positive rate due to routing dynamics affecting anycast detection. To overcome this issue, the authors proposed a two-step system. However, they did not provide a single integrated solution to realize such measurements. Furthermore, neither MAnycast2 nor iGreedy was designed for daily anycast censuses. We wanted to build a system that: i) combines anycast- and latency-based GCD measurements; ii) is accurate and precise; and iii) performs a responsible measurement in terms of probing budget. We also want to release a continual daily census for the community and release the toolchain for others to build on. To achieve these goals we set the following requirements:
@@
    Accurate - MAnycast2 suffers from false positives and false negatives. LACeS must minimise such cases, and must convey confidence in results through independently listing the classification for the anycast-based and GCD approach.
@@
    Precise - Anycast detection errors between subsequent censuses must be minimized to support longitudinal measurements. In other words: LACeS must be precise over time.
@@
    Responsible - LACeS must perform measurements with low impact on the platform it runs on and the Internet as a whole. LACeS must support low probing rates without impacting accuracy and precision.
@@
    Multi-protocol - iGreedy and MAnycast2 only support ICMPv4. LACeS must also support IPv6, transport layer protocols (TCP, UDP) and service-aware probing.
@@
    Performant - MAnycast2 had poor scalability and iGreedy cannot perform a comprehensive IPv4 census in a single day. LACeS must complete both a hitlist-based IPv4 census and IPv6 census for multiple protocols in a single day.
@@
    Reproducible - Independent parties, that may have anycast deployments of various sizes and limited resource availability, must be able to run LACeS.
@@ 4. Design and Measurement Setup
We begin with a set of responsive addresses across the Internet (§4.1). The LACeS probing system consists of two components: one based on MAnycast24.2) that detects candidate anycast prefixes, and one based on iGreedy that infers if those prefixes are anycast, and the location of sites for anycast prefixes (§4.3).
4.1. Census Input
We use ISI’s IPv4 hitlist (Fan and Heidemann, 2010), which ranks ping-responsive addresses per /24 prefix, to select ping targets. For DNS we merge the ISI hitlist with a set of authoritative name server IPs obtained from the OpenINTEL project (van Rijswijk-Deij et al., 2016). We prefer name server IPs as representative addresses for a specific /24 to maximize chances of probing an active DNS server. Next, we created a TCP hitlist based on Zmap (Durumeric et al., 2013) scans targeting the entire routable IPv4 address space. These hitlists cover 6.0 M /24s, with a union of 4.3 M responsive (3.8 M ICMP, 2.9 M TCP, 0.2 M DNS) across 73k ASes.
For IPv6 we use the public hitlist from TU Munich (TUM) (Gasser et al., 2018) combined with AAAA DNS records obtained from OpenINTEL. We consider only the first /48 of large aliased prefixes considered to be allocated to the same service, to maintain reasonable probing burden and times. This leaves 6.2 M /48s, with a union of 5.6 M responsive (5.3 M ICMP, 4.3 M TCP, 26k DNS) across 21k ASes. Due to the longitudinal aspect of LACeS we update hitlists quarterly, in sync with ISI’s hitlists (Fan and Heidemann, 2010), to maintain coverage.
We probe at a /24-IPv4-prefix and /48-IPv6-prefix granularity, targeting a single address per prefix, as these are likely to be the smallest prefix size to be propagated by BGP. This reduces the probing burden and makes daily censuses feasible. Whilst this assumption works well in most cases, Schomp et al. showed that there may be rare cases of more specific partitioning (Schomp and Al-Dalky, 2020). For example, hypergiants, such as Microsoft, announcing a /24 prefix at their PoPs may assign part of a prefix to a replicated service, whilst utilizing all other addresses in that prefix as unicast for individual servers within their private backbone, we refer to this practice as partial anycast. While we acknowledge this limitation, we continue to probe at a /24 prefix granularity to keep our scanning rate low for responsible scanning. We address and quantify this limitation in §5.7.
4.2. Detecting Candidate Anycast Prefixes
Figure 3. Overview of the measurement pipeline.